### v3

@@ Appendix 0.B Safe sequences in arc-path covers
We work on $s$-$t$ DAGs with node set $N$ and arc set $A$ and allow parallel arcs, so that in fact are working on directed acyclic multigraphs. For $X\subseteq A$ we denote a sequence of arcs $X$ through arcs $a_{1},a_{2},\dots,a_{k}$ as $X=a_{1},\dots,a_{k}$ if each arc in the sequence reaches the one after it; a path is now denoted as a contiguous sequence of arcs. If $Y$ is a sequence of arcs and each arc of $X$ is contained in $Y$ we say that $X$ is a subsequence of $Y$, or that $X$ is contained in $Y$. Two paths are distinct if they differ in at least one arc, and two path cover are distinct if they differ in at least one path. A sequence of arcs is safe for a DAG $G$ if it is a subsequence of some $s$-$t$ path in every path cover of $G$.
We say that $xy$ is a $u$-$v$ bridge if there is no $u$-$v$ path in $G$ without $xy$; equivalently, all $u$-$v$ paths in $G$ contain the arc $xy$. The concepts introduced for dominators in Section˜2 can be defined for arcs: given an arc $uv$, we say that xyxy $s$-dominates uvuv iff every $s$-$u$ path contains $xy$, and zwzw $t$-dominates uvuv iff every $v$-$t$ path contains $zw$. We can also refine the notion of extension for arcs by saying that $\mathsf{extension}(uv)$ consists of the sequence of $s$-$u$ bridges concatenated with $uv$, followed by the sequence of $v$-$t$ bridges.
A unitary path with respect to arc-path covers is a path where each node that is not the first nor the last has indegree and outdegree equal to one. The arc-compressed graph of $G$ is a graph where each maximal unitary path is replaced by an arc from the first node to the last node of the path. Note that in an arc-compressed graph, no node has both indegree and outdegree equal to one. Note also that we might introduce parallel arcs during compression if a transitive arc exists from the first to the last node of the unitary path, so indeed it is important that we allow multigraphs to begin with. Arc-compressing a graph can be done in linear time, similarly to point 1 of Lemma˜1. The line graph of a directed graph $G=(V,E)$, denoted by $L(G)$, is a graph where each node corresponds to an arc of $G$ and where there is an arc from node $uv$ to node $xy$ if $v=x$; essentially, arcs of $L(G)$ encode paths of two arcs in $G$.
@@
    There is an arc $uv\in A$ such that every $s$-$t$ path covering $uv$ contains $X$.
@@
$(\Leftarrow)$ If the unique in-neighbor of $v$ is $u$ and there exists an arc $vw$, then $uv$ is the immediate dominator of $vw$. In fact, $uv$ is the immediate dominator of every outgoing arc from $v$.
$(\Rightarrow)$ Suppose for a contradiction that $d_{v}^{+}<1$ or $d_{v}^{-}\neq 1$. If $d_{v}^{+}=0$ then $v=t$ and so $uv$ is not a dominator, so suppose $v$ has at least one out-neighbor. If $v$ has an in-neighbor $z\neq u$ then there is an $s$-$z$ path avoiding $uv$, for otherwise there would be a cycle in $G$ through $v$. Therefore $uv$ does not dominate any of the outgoing arcs from $v$, and thus it does not dominate any of the arcs it reaches: to any path containing $uv$ we change the prefix of the path until $v$ with an $s$-$v$ path through $z$ avoiding $uv$. $\square$
@@
Let $uv$ be an $s$-dominator. By Lemma˜5 we have $d^{-}_{v}=1$, and so $uv$ immediately dominates each outgoing arc from $v$. Since $G$ is compressed we must have $d_{u}^{+}\geq 2$, proving the statement. $\square$
@@
Let $G=(N,A)$ be an $s$-$t$ DAG and let $uv,xy\in A$ be arcs. If $uv$ immediately $s$-dominates xyxy, then the immediate $t$-dominator of uvuv $t$-dominates xyxy.
@@
Let $a$ denote the node corresponding to $uv$ in $L(G)$ and analogously $b$ to $xy$. Apply Lemma˜3 to $L(G)$, $a$, $b$, and $k=1$. $\square$
In a compressed graph, whenever $uv$ immediately $s$-dominates xyxy and $xy$ immediately $t$-dominates uvuv, then $v\neq x$. Otherwise, by Lemma˜5, we would have $d^{-}_{v}=1$ because $uv$ is an $s$-dominator and $d^{+}_{v}=1$ because $xy$ is a $t$-dominator, which contradicts the assumption that the graph is compressed.
@@
Let $G=(N,A)$ be an arc-compressed $s$-$t$ DAG and let $uv\in A$ be an arc. Then $\mathsf{extension}(uv)$ is a maximal safe sequence of arcs if and only if $uv$ is a leaf in both the $s$- and $t$-dominator trees of G.
@@
$(\Leftarrow)$ If $uv$ is a leaf in both dominator trees then the sequence that $uv$ identifies via paths to the roots of the dominator trees is maximal: no other arc can be added to this sequence by the equivalence of safe sequences and paths in the dominator trees given by Theorem˜0.B.1.
$(\Rightarrow)$ Suppose that $uv$ is not a leaf in the dominator tree of $s$. We apply Lemma˜6 to $G$ to get two arcs $s$-dominated by $uv$, xyxy and $zw$. It suffices to show that the lowest common ancestor in the $t$-dominator tree between $uv$ and $xy$ or between $uv$ and $zw$ is the immediate $t$-dominator of $uv$. By Lemma˜7, we can assume that $xy$ is not the immediate $t$-dominator of $uv$. Let $a$ denote the lowest common ancestor in the $t$-dominator tree between $uv$ and $xy$. If $a$ is not the immediate $t$-dominator of $uv$ then it is an ancestor of it. Consider an xyxy-$t$ path avoiding the immediate dominator of $uv$ and consider a uvuv-xyxy path avoiding the immediate $t$-dominator of $uv$; these paths together form a uvuv-$t$ path avoiding the immediate $t$-dominator of $uv$, a contradiction. $\square$
@@ Appendix 0.D Applying safe sequences to simplify path-covering ILPs
For part (i), for each solution path $i$ that needs to be encoded ($i\in\{1,\dots,k\}$), both [17, 48] introduce a binary variable $x_{uvi}$ associated to each arc $uv$ of the DAG, with the interpretation that $x_{uvi}$ equals 1 if and only if the arc $uv$ belongs to the $i$-th solution path. Since the input graph is a DAG, this can be guaranteed by stating that the sum of all $x_{sui}$ variables on arcs $su$ exiting the source $s$ is exactly 1, and for all nodes $v$ (different from $s$ and $t$) the sum of the binary variables $x_{uvi}$ of arcs $uv$ entering $v$ is equal to the sum of the binary variables $x_{vwi}$ of arcs $vw$ exiting $v$, see [17].
@@
In this paper, we apply safe sequences to fix some $x_{uvi}$ variables to $1$, thus reducing the search space of the ILP solver, using the approach of Grigorjew et al. [25]. Let $X_{1},\dots,X_{t}$ be sequences in the input DAG such that: (a) each $X_{i}$, $i\in\{1,\dots,t\}$, must appear in some solution path of the ILP, and (b) there exists no path in the DAG containing two distinct $X_{i}$ and $X_{j}$. In the case that every $X_{i}$ is a path, Grigorjew et al. [25] noticed that if the two mentioned properties hold, then $X_{1},\dots,X_{t}$ must appear in $t$ distinct paths of among the $k$ solution paths of the ILP. As such, one can, without loss of generality, assign each $X_{i}$ to the $i$-th solution path. That is, for each $i\in\{1,\dots,t\}$, we set $x_{uvi}=1$ for each arc $uv$ of $X_{i}$.
Finally, if it holds that the ILP solution paths cover all arcs of the graph or all arcs in the set $C$ with respect to which we compute safe sequences, then we can use safe sequences in the above procedure to ensure that the ILP still has the same set of optimal solutions. To find suitable sets of sequences, and to have a large number of binary variables set to 1 in this manner, we follow the approach of [25]. Namely, for every arc $uv$, we record the longest safe sequence containing $uv$, and use its length as the weight of $uv$. Then, we find a maximum-weight antichain of arcs in the DAG, namely a set of arcs $a_{i},\dots,a_{t}$ such that from no $a_{i}$ we can reach some other $a_{j}$, $i,j\in\{1,\dots,t\}$, and its total weight is maximum among all antichains, using the min-flow reduction of [40]. Then, for each $a_{i}$ in the maximum-weight antichain, we take the recorded longest safe sequence $X_{i}$. To see that there can be no path in the DAG containing distinct $X_{i}$ and $X_{j}$, notice that such path would contain both the corresponding arcs $a_{i}$ and $a_{j}$, which contradicts the fact that $a_{i}$ and $a_{j}$ belong to an antichain.
@@
Here we describe two path-covering ILPs that we chose to experiment with. For both, we assume that for every arc $uv$ of the graph we have an associated positive weight $w(uv)$. We also assume part (i) discussed above, namely, the $x_{uvi}$ binary variables, and the constraints discussed above to ensure they induce paths. The overall goal is to find the “best” $k$ such paths, and their associated weights $w_{1},\dots,w_{k}$.

### v4

@@ Appendix 0.B Safe sequences in arc-path covers
We work on $s$-$t$ DAGs with node set NN and arc set AA and allow parallel arcs, so that in fact are working on directed acyclic multigraphs. For $X\subseteq A$ we denote a sequence of arcs $X$ through arcs $a_{1},a_{2},\dots,a_{k}$ as $X=a_{1},\dots,a_{k}$ if each arc in the sequence reaches the one after it; a path is now denoted as a contiguous sequence of arcs. If $Y$ is a sequence of arcs and each arc of $X$ is contained in $Y$ we say that $X$ is a subsequence of $Y$, or that $X$ is contained in $Y$. Two paths are distinct if they differ in at least one arc, and two path cover are distinct if they differ in at least one path. A sequence of arcs is safe for a DAG $G$ if it is a subsequence of some $s$-$t$ path in every path cover of $G$.
We say that xyxy is a $u$-$v$ bridge if there is no $u$-$v$ path in $G$ without xyxy; equivalently, all $u$-$v$ paths in $G$ contain the arc xyxy. The concepts introduced for dominators in Section˜2 can be defined for arcs: given an arc uvuv, we say that xyxy $s$-dominates uvuv iff every $s$-$u$ path contains xyxy, and zwzw $t$-dominates uvuv iff every $v$-$t$ path contains zwzw. We can also refine the notion of extension for arcs by saying that $\mathsf{extension}(uv)$ consists of the sequence of $s$-$u$ bridges concatenated with uvuv, followed by the sequence of $v$-$t$ bridges.
A unitary path with respect to arc-path covers is a path where each node that is not the first nor the last has indegree and outdegree equal to one. The arc-compressed graph of $G$ is a graph where each maximal unitary path is replaced by an arc from the first node to the last node of the path. Note that in an arc-compressed graph, no node has both indegree and outdegree equal to one. Note also that we might introduce parallel arcs during compression if a transitive arc exists from the first to the last node of the unitary path, so indeed it is important that we allow multigraphs to begin with. Arc-compressing a graph can be done in linear time, similarly to point 1 of Lemma˜1. The line graph of a directed graph $G=(V,E)$, denoted by $L(G)$, is a graph where each node corresponds to an arc of $G$ and where there is an arc from node uvuv to node xyxy if $v=x$; essentially, arcs of $L(G)$ encode paths of two arcs in $G$.
@@
    There is an arc $uv\in A$ such that every $s$-$t$ path covering uvuv contains $X$.
@@
$(\Leftarrow)$ If the unique in-neighbor of $v$ is $u$ and there exists an arc vwvw, then uvuv is the immediate dominator of vwvw. In fact, uvuv is the immediate dominator of every outgoing arc from $v$.
$(\Rightarrow)$ Suppose for a contradiction that $d_{v}^{+}<1$ or $d_{v}^{-}\neq 1$. If $d_{v}^{+}=0$ then $v=t$ and so uvuv is not a dominator, so suppose $v$ has at least one out-neighbor. If $v$ has an in-neighbor $z\neq u$ then there is an $s$-$z$ path avoiding uvuv, for otherwise there would be a cycle in $G$ through $v$. Therefore uvuv does not dominate any of the outgoing arcs from $v$, and thus it does not dominate any of the arcs it reaches: to any path containing uvuv we change the prefix of the path until $v$ with an $s$-$v$ path through $z$ avoiding uvuv. $\square$
@@
Let uvuv be an $s$-dominator. By Lemma˜5 we have $d^{-}_{v}=1$, and so uvuv immediately dominates each outgoing arc from $v$. Since $G$ is compressed we must have $d_{u}^{+}\geq 2$, proving the statement. $\square$
@@
Let $G=(N,A)$ be an $s$-$t$ DAG and let $uv,xy\in A$ be arcs. If uvuv immediately $s$-dominates xyxy, then the immediate $t$-dominator of uvuv $t$-dominates xyxy.
@@
Let $a$ denote the node corresponding to uvuv in $L(G)$ and analogously $b$ to xyxy. Apply Lemma˜3 to $L(G)$, $a$, $b$, and $k=1$. $\square$
In a compressed graph, whenever uvuv immediately $s$-dominates xyxy and xyxy immediately $t$-dominates uvuv, then $v\neq x$. Otherwise, by Lemma˜5, we would have $d^{-}_{v}=1$ because uvuv is an $s$-dominator and $d^{+}_{v}=1$ because xyxy is a $t$-dominator, which contradicts the assumption that the graph is compressed.
@@
Let $G=(N,A)$ be an arc-compressed $s$-$t$ DAG and let $uv\in A$ be an arc. Then $\mathsf{extension}(uv)$ is a maximal safe sequence of arcs if and only if uvuv is a leaf in both the $s$- and $t$-dominator trees of G.
@@
$(\Leftarrow)$ If uvuv is a leaf in both dominator trees then the sequence that uvuv identifies via paths to the roots of the dominator trees is maximal: no other arc can be added to this sequence by the equivalence of safe sequences and paths in the dominator trees given by Theorem˜0.B.1.
$(\Rightarrow)$ Suppose that uvuv is not a leaf in the dominator tree of $s$. We apply Lemma˜6 to $G$ to get two arcs $s$-dominated by uvuv, xyxy and zwzw. It suffices to show that the lowest common ancestor in the $t$-dominator tree between uvuv and xyxy or between uvuv and zwzw is the immediate $t$-dominator of uvuv. By Lemma˜7, we can assume that xyxy is not the immediate $t$-dominator of uvuv. Let $a$ denote the lowest common ancestor in the $t$-dominator tree between uvuv and xyxy. If $a$ is not the immediate $t$-dominator of uvuv then it is an ancestor of it. Consider an xyxy-$t$ path avoiding the immediate dominator of uvuv and consider a uvuv-xyxy path avoiding the immediate $t$-dominator of uvuv; these paths together form a uvuv-$t$ path avoiding the immediate $t$-dominator of uvuv, a contradiction. $\square$
@@ Appendix 0.D Applying safe sequences to simplify path-covering ILPs
For part (i), for each solution path $i$ that needs to be encoded ($i\in\{1,\dots,k\}$), both [17, 48] introduce a binary variable $x_{uvi}$ associated to each arc uvuv of the DAG, with the interpretation that $x_{uvi}$ equals 1 if and only if the arc uvuv belongs to the $i$-th solution path. Since the input graph is a DAG, this can be guaranteed by stating that the sum of all $x_{sui}$ variables on arcs susu exiting the source $s$ is exactly 1, and for all nodes $v$ (different from $s$ and $t$) the sum of the binary variables $x_{uvi}$ of arcs uvuv entering $v$ is equal to the sum of the binary variables $x_{vwi}$ of arcs vwvw exiting $v$, see [17].
@@
In this paper, we apply safe sequences to fix some $x_{uvi}$ variables to 11, thus reducing the search space of the ILP solver, using the approach of Grigorjew et al. [25]. Let $X_{1},\dots,X_{t}$ be sequences in the input DAG such that: (a) each $X_{i}$, $i\in\{1,\dots,t\}$, must appear in some solution path of the ILP, and (b) there exists no path in the DAG containing two distinct $X_{i}$ and $X_{j}$. In the case that every $X_{i}$ is a path, Grigorjew et al. [25] noticed that if the two mentioned properties hold, then $X_{1},\dots,X_{t}$ must appear in $t$ distinct paths of among the $k$ solution paths of the ILP. As such, one can, without loss of generality, assign each $X_{i}$ to the $i$-th solution path. That is, for each $i\in\{1,\dots,t\}$, we set $x_{uvi}=1$ for each arc uvuv of $X_{i}$.
Finally, if it holds that the ILP solution paths cover all arcs of the graph or all arcs in the set $C$ with respect to which we compute safe sequences, then we can use safe sequences in the above procedure to ensure that the ILP still has the same set of optimal solutions. To find suitable sets of sequences, and to have a large number of binary variables set to 1 in this manner, we follow the approach of [25]. Namely, for every arc uvuv, we record the longest safe sequence containing uvuv, and use its length as the weight of uvuv. Then, we find a maximum-weight antichain of arcs in the DAG, namely a set of arcs $a_{i},\dots,a_{t}$ such that from no $a_{i}$ we can reach some other $a_{j}$, $i,j\in\{1,\dots,t\}$, and its total weight is maximum among all antichains, using the min-flow reduction of [40]. Then, for each $a_{i}$ in the maximum-weight antichain, we take the recorded longest safe sequence $X_{i}$. To see that there can be no path in the DAG containing distinct $X_{i}$ and $X_{j}$, notice that such path would contain both the corresponding arcs $a_{i}$ and $a_{j}$, which contradicts the fact that $a_{i}$ and $a_{j}$ belong to an antichain.
@@
Here we describe two path-covering ILPs that we chose to experiment with. For both, we assume that for every arc uvuv of the graph we have an associated positive weight $w(uv)$. We also assume part (i) discussed above, namely, the $x_{uvi}$ binary variables, and the constraints discussed above to ensure they induce paths. The overall goal is to find the “best” $k$ such paths, and their associated weights $w_{1},\dots,w_{k}$.